A GENERAL RATIO ESTIMATOR AND ITS APPLICATION IN MODEL BASED INFERENCE

Z. OUYANG^{1}, J. N. SRIVASTAVA^{2} AND H. T. SCHREUDER^{1}

1240 W. Prospect St., Rocky Mountain Forest Experiment Station, Fort Collins, CO 80521, U.S.A. 2Department of Statistics, Colorado State University, Fort Collins, CO 80523, U.S.A.

(Received October 29, 1990; revised April 16, 1992)

Abstract. A general ratio estimator of a population total is proposed as an approximation to the estimator introduced by Srivastava (1985, Bull. Internat. Statist. Inst., 51(10.3), 1–16). This estimator incorporates additional information gathered during the survey in a new way. Statistical properties of the general ratio estimator are given and its relationship to the estimator proposed by Srivastava is explored. A special kind of general ratio estimator is suggested and it turns out to be very efficient in a simulation study when compared to several other commonly used estimators.

Key words and phrases: Finite population sampling, sample weight function, linear model, regression estimator, the Horvitz-Thompson estimator.